OpenStudy (anonymous):

If m and n are 2 numbers which belong to {1,2,3...100} then find Probability that 7^m + 7^n is divisible by 5.

OpenStudy (raden):

there are 4 pairs (m,.n) which satisfies : (1,3),(3,1),(2,4),(4,2) while all pairs 10P2 can you figure out the answer then ...

OpenStudy (anonymous):

sorry missed a 0 :P

OpenStudy (raden):

what, 100 ? bahhhh :)

OpenStudy (anonymous):

sorry :|

OpenStudy (raden):

see the result of exponents of 7 (just the units) 7^1 = 7 7^2 = ...9 7^3 = ...3 7^4 = ...1 7^5 = ...7 7^6 = ...9 7^7 = ...3 7^8 = ...1 ...... ...... and so on as the total numberof outcames = 100P2 but still wonder how to get number of events, if i used trial too longggg if m = 1, then n = {3,7,11,15,19,23,27,...,99}. there are 25 events if m = 5, then n = {3,7,11,15,19,23,27,...,99}. there are 25 events too ..... --"

OpenStudy (anonymous):

hmm..

ganeshie8 (ganeshie8):

continuing rad, im getting :- 2[25x25] + 2[25x25] probability = ---------------------- 100x100 = 1/4